Sorry Jon!
It is
x + 7y = -39
-3x - 2y = 22
Substitution method ,I can't get it to work out.

Answer by longjonsilver(2297)

let 2 eqns be eqn1 and eqn2, respectively.

in eqn1, x = -39 - 7y. So put this (ie substitute it) into the second eqn, which becomes:

-3(-39-7y) - 2y = 22 --> expand the bracket and simply it all:
117+21y - 2y = 22
19y = 22-117
19y = -95
y = -95/19
y = -5.

So now put this into either eqn1 or eqn2..pick the easiest looking one!... eqn2

This becomes -3x -2(-5) = 22
-3x + 10 = 22
-3x = 22-10
-3x = 12
x = -12/3
x = -4

Now but both answers into the other eqn and see if you do get -39. If you do, you KNOW the answer is correct!

-4 + 7(-5)
-4 - 35
-39 --> CORRECT!!

jon.